Suppose another bank offers you this deal? You can to invest 10,000 into their 3-year CD. But they are offering a yearly interest rate of 6%, compounded continuously. How much would the investment mature to in 3 years?
1
Expert's answer
2015-10-12T00:00:45-0400
If another bank offers you this deal and you can invest 10,000 into their 3-year CD, but they are offering a yearly interest rate of 6%, compounded continuously. In this case the investment will mature in 3 years to A = P*e^rt = 10,000*e^0.06*3 = 11,972.17.